OpenStudy (anonymous):

If 0.98 moles of BaCl2 are dissolved in enough water to make a 1.2-liter solution, what is the resulting molarity

OpenStudy (asib1214):

0.98/1.2= 0.816M

OpenStudy (abhisar):

Hello @dreamchaser22 ! \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

\(\sf \large Molarity=\frac{No.~of~moles~of~solute}{Volume~of~solution~in~litre}\)

OpenStudy (abhisar):

Here, moles of solute=0.98 Volume of solution=1.2 L \(\sf \large Molarity=\frac{0.98}{1.2}\)